OpenStudy (anonymous):

Solve for x: -6 + 9 |x + 3| = 12

OpenStudy (camerondoherty):

First add -6 + 9

OpenStudy (camerondoherty):

which equals 3

OpenStudy (camerondoherty):

then subtract 3 from both sides of the equation

OpenStudy (camerondoherty):

now u have: |x+3|=9

OpenStudy (camerondoherty):

What do you think we do next?

OpenStudy (juanpablojr):

It would be so much shorter to understand here http://bit.ly/1q1Jg5g

OpenStudy (camerondoherty):

lol

OpenStudy (anonymous):

@camerondoherty why wouldnt you add 6 to both sides then divide by 9?

OpenStudy (anonymous):

=69

OpenStudy (camerondoherty):

because you always combine like terms first

OpenStudy (camerondoherty):

-6 and 9 are like terms

OpenStudy (anonymous):

Its a multiple choice question, a. {-7, 1 } b. {-12,6 } c. {4, 7} d. {-5, -1}

OpenStudy (camerondoherty):

Probably D

OpenStudy (anonymous):

she wants the D

OpenStudy (anonymous):

LMAO

OpenStudy (anonymous):

amirite?

OpenStudy (anonymous):

that was rude

OpenStudy (camerondoherty):

^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

but cam D is if you add 6 then divide thats how i was doing it then checked my work with an online calculator and it did it the same way the way you were doing it combining terms you would get B

OpenStudy (anonymous):

kio is right, the 9 is multiplied

OpenStudy (anonymous):

no but it is defiantly D

OpenStudy (camerondoherty):

oh :/ I guess i messed up *sigh* i always mess up lol

OpenStudy (anonymous):

lol medal for trying :)

OpenStudy (camerondoherty):

thx c:

OpenStudy (anonymous):

\[\large -6 + 9|x+3| = 12\] \[\large 9|x+3| = 18\] \[\large |x+3| = 2\] \[\large x+3 = -2 ~~~~~ x + 3 = 2 \] \[\large \text{solve both now.} \]

OpenStudy (anonymous):

http://symbolab.com/solver/absolute-equation-calculator/-6%2B9%7Cx%2B3%7C%3D12/?origin=button this site is useful to solve absolute values and shows step by step if we arent explaining well @Bullet145

OpenStudy (anonymous):

@camerondoherty cant reply to email you have it to where you can only email ppl you are fanned to

OpenStudy (anonymous):

mathway.com is also very helpful

OpenStudy (anonymous):

i think the answer is D

OpenStudy (cj49):

x=-5 or x=-1